1. Compute the force of attraction between two planets (Planet A and Planet B). The mass of Planet A is 12.34 x 1024 kg and Planet B is 364.25 x 1020 kg and the distance between them is 788,354 km. 



1
Expert's answer
2022-06-07T17:37:42-0400
"F=G\frac{m_1m_2}{r^2}"

"F=6.67*10^{-11}*\frac{12.34*10^{24}*364.25*10^{20}}{(788.354*10^6)^2}\\
=4.82*10^{19}\:\rm N"